OpenStudy (anonymous):

Mary works in a factory where she is paid $5 per hour up to 40 hours. If she works over 40 hours, she is paid twice her usual wage. Write a linear equation to model what would be used to determine her pay for the week.

hero (hero):

y = 5*40 + 10x x = overtime hours

OpenStudy (anonymous):

what does the 5*40 mean?

hero (hero):

5*40 = 5 x 40